009 10.0 points The driver of a car traveling at 36.3 m/s ap- plies the brakes and undergoes a constant deceleration of 1.28 m/s How many revolutions does each tire make before the car comes to a stop, assuming that the car does not skid and that the tires have radii of 0.27 m? Answer in units of rev
Answer #1

Using kinematics, we have

v = u+ at

0 = 36.3 + (-1.28)t

t = 28.35 sec

Now, we will find the distance covered in this time.

d = ut + 1/2at2

d = 36.3*28.35 + 1/2 (-1.28)28.352

d = 1029.45 - 514.4

d = 515.1 m

Now, circumference of tire is

C = 2*pi*r = 1.69646 m

so,

number of revolutions = 515.1 / 1.69646

number of revolutions = 303.6

so, number of revs = 303